OpenStudy (anonymous):

HELP PLEASE! The graph below shows the revenue earned by a company in its first four quarters of operation. A line of best fit has been drawn on the graph. A first−quadrant grid with 'Revenue (in 1000 dollars)' as they y−axis label and 'Number of Quarters' as the x-axis label. The y axis ranges from 0 to 55 with an interval of 5 units. The x axis ranges from 0 to 7 with an interval of 1 unit. There are points at the following locations: (1, 1)(2, 25)(3, 30)(4, 40). A straight line is drawn through the points and passes through (1, 1)(3, 30) and (4, 40).

OpenStudy (anonymous):

Based on the linear model, how much revenue, in thousands of dollars, should the company earn in its 8th quarter of operation? Type your answer in the box below.

OpenStudy (anonymous):

@jim_thompson5910 can you help?

OpenStudy (anonymous):

Is the coordinates (1 , 1) or (1 , 10)

OpenStudy (anonymous):

I am assuming it's (1 , 10) or else it's not a straight line anymore

OpenStudy (anonymous):

its 1,10

OpenStudy (anonymous):

Find the equation of the line, using y = y1 + m(x - x1) where m is the gradient and (x1 , y1) is a coordinates of a point on the line. Gradient = (30 - 10) / (3 - 1) = 20 / 2 = 10 Using (1 , 10): y = 10 + 10(x - 1) = 10 + 10x - 10 = 10x x = 8 y = 10(8) = 80

OpenStudy (anonymous):

so its 80,000?